# Fading in the Flow: Suppression of cold gas growth in expanding galactic outflows

### v2

@@ 4.3 Caveats and Future Directions
The Chevalier & Clegg 1985 (CC85) model of an adiabatically expanding outflow makes dramatic simplifications. Observations show that the thermodynamic state of the outflowing gas can be significantly different from a steady CC85 wind (Xu et al. 2023). For instance, Lopez et al. 2020 observe much shallower density and temperature profiles (also in free electron density profile observed by Deanne Fisher; private communication) for the X-ray bright hot gas beyond $1\ \rm kpc$ in M82 compared to the CC85 model. Such deviations can be generic and wind profiles can potentially be non-steady in various starburst-driven outflows. Our work highlights that the expansion of the wind bears non-trivial implications for the survival/growth of cold clouds in outflows. [^10] For instance, in an isothermally expanding wind, we expect a significantly reduced cloud mass growth than in an adiabatically expanding wind. From our analytic considerations, we expect $m_{\rm cl}\propto d_{\rm cl}^{1/4}$ in an isothermal wind (not validated against any simulations carried out in this paper). This is not surprising, and depending on the background profile, clouds behave significantly differently in how they mix, shred, or grow (cf. also Tan et al. 2023 for clouds infalling in a stratified medium).
@@
An extensive study of the behavior of cold clouds in diverse astrophysically relevant environments under a unified framework has yet to be explored. Furthermore, the assumptions that go into the CC85 model, such as spherical symmetry and continuous energy and mass injection are extremely simplistic. Supernovae can be clustered or spread out in the galactic disk, and the injection of mass and energy by them can be in the form of bursty episodes (Schneider et al. 2020; Schneider & Mao 2024; Mohapatra et al. 2025), for which the outflow does not attain a smooth CC85 profile (Sharma et al. 2014). All of these effects can add additional complications, including driving a significant amount of turbulence in the wind ( Ghosh et al.; in prep) and transport of metals (Sharda et al. 2024) that can potentially alter how the individual clouds or their population evolve in outflows. Therefore, different models of supernovae feedback (Warren et al. 2025) and turbulence give rise to a great deal of spatial and temporal variations in the background conditions (Mohapatra & Sharma 2019; Mohapatra et al. 2020) that different cold clouds encounter. These must be carefully taken into account before robust comparison can be made with observations and simulations (both large-scale outflow simulations of isolated galaxies; Schneider & Robertson 2017; Schneider et al. 2018, 2020; Kim & Ostriker 2018; Vijayan et al. 2018, 2020; Steinwandel et al. 2024; Wang et al. 2024; Li et al. 2025; Vijayan et al. 2025 or cosmological simulations of galaxy formation; Davé et al. 2020; Ramesh et al. 2024) of cold clouds and their host galactic winds. Although not directly related to the idealized simulations discussed here, similar caveats must be considered while testing the applicability of different sub-grid models of cold gas in cosmological simulations (Weinberger & Hernquist 2023; Smith et al. 2024a, b; Bennett et al. 2024; Butsky et al. 2024; Das et al. 2024) under various kinematic and thermodynamic conditions.
@@
In addition, the effects of thermal conduction and magnetic fields can be significant in determining the small-scale structure of cold clouds. Several recent works have demonstrated that magnetic fields suppress mixing-induced hydrodynamic instabilities, reducing the rate at which the cold mass grows in cloud-crushing simulations (McCourt et al. 2015; Kooij et al. 2021) – a picture also supported by recent idealized simulations of radiative turbulent mixing layers with magnetic fields by Das & Gronke 2024. Moreover, magnetic fields give rise to additional drag on the cloud due to draping by magnetic fields (Dursi & Pfrommer 2008) which is very sensitive to not just the magnetic field strength but more importantly to the orientation of the magnetic fields (Hidalgo-Pineda et al. 2024; Kaul et al. 2025) with respect to the direction of motion of the clouds or the direction of spatial gradients in the background medium. Strength and orientation of magnetic fields might also be responsible for the generation of cold filamentary structures observed in galactic outflows (like M82) that bear striking similarity with solar prominences. In systems similar to cloud-crushing, magnetic fields have also been found to modulate the properties of ram-pressure stripped tails in jellyfish galaxies (Müller et al. 2021). Therefore, magnetic fields and their associated effects can give rise to slow movement of long-lived clouds, leading to higher cold-mass seeding at any position in the bulk of the outflow, albeit at a lower rate. Furthermore, a distribution of multiple clouds (common in the CGM; cf. Liang & Remming 2020; Singh Bisht et al. 2024; Hummels et al. 2024; Tan & Fielding 2024) can interact with each other (Nguyen et al. 2024) and can also change the bulk properties of the wind by inducing precipitation (owing to thermal instability; Thompson et al. 2016; Schneider et al. 2018; Dannen et al. 2020; Lochhaas et al. 2021). Such effects can give rise to multiphase outflows with significantly different kinematic and thermodynamic properties, not considered in our work. Careful consideration of such effects and their relative importance and diverse outflow environments should be considered in the future. This would enable simulations to make realistic predictions of cloud properties in different regimes and bridge the discrepancies between predictions from small-scale simulations and results of large-scale simulations of individual galaxies or cosmological simulations.
@@ 5 Summary & Conclusion
The formation and survival of cold clouds in galactic outflows are crucial for explaining the ubiquity of cold gas in the CGM. Observations suggest that cold clouds are accelerated farther from the galactic disk by powerful bursts of supernovae. The outflowing gas is routinely found to have an expansion in the wind away from the center of the galactic disk. In this work, we simulate the evolution of cold clouds seeded in an expanding wind, which is modeled by a steady-state Chevalier & Clegg 1985 (CC85) profile. The adiabatically expanding outflow shreds, stretches, and mixes the cold clouds seeded in the wind. In general, we find that
@@
    Compared to the indefinite growth of cold mass seen in ’vanilla’ cloud crushing simulations (Brüggen & Scannapieco 2016; Armillotta et al. 2017; Gronke & Oh 2018; Li et al. 2020; Sparre et al. 2020; Kanjilal et al. 2021; Abruzzo et al. 2023), clouds with the same initial condition but exposed to a steady adiabatically expanding (CC85) wind, instead of a uniform wind, undergo significantly lower cold mass growth (can be lower by almost an order of magnitude; see Fig. 5).
@@
    We find that the expansion associated with the CC85 wind results in clouds embedded in them to spread in directions orthogonal to the wind. Once entrained, the cloud expansion is proportional to the distance travelled by the cloud $d_{\rm cl}$ (see Fig. 9), and the cold gas mass $m_{\rm cl}$ grows proportional to the square of the distance they travel, i.e., $m_{\rm cl}\propto d_{\rm cl}^{2}$ (for entrained clouds; see Fig. 10). This also explains the suppression of cold mass growth compared to the uniform wind ‘vanilla’ cloud-crushing, where $m_{\rm cl}\propto d_{\rm cl}^{3}$ – a significantly steeper cold mass growth with distance.
@@
    We compare our simulation results with the prediction of the phenomenological model of single-cloud evolution in galactic outflows proposed by Fielding & Bryan 2022 (FB22; Section 3 in their paper). Upon comparing cloud mass growth with time with our simulations using the parameters chosen for their model (Section 3.2 in their paper), we are unable to reproduce the suppression of cold mass growth in a CC85 wind compared to their ‘vanilla’ cloud-crushing counterparts (uniform wind) for all our simulation parameters. The FB22 model is very sensitive to the fractal dimension of the radiatively cooling turbulent boundary layer, which can be tweaked to match our simulation result (see Fig. 12). This highlights that the treatment of an individual cloud evolution in FB22 (Section 3.1 in their paper) is generic and can potentially be improved to match simulation results in future. However, this is not robust across different simulation parameters and further work is needed towards developing a sound physical interpretation of any specific choice of single-cloud evolution model. FB22 also developed a general framework for the profile of a multiphase outflow (equations in Fig. 3 & Eq. 17 in FB22) with different energy and mass source terms in the wind. These source terms can come from any model of single cloud evolution in galactic outflows. In particular, we introduce source terms to account for the cloud-wind interaction from the phenomenological cloud-crushing model that we discussed earlier. Using the general multiphase wind framework from FB22, we find that changes in the behavior of individual clouds in the wind lead to substantial changes in the global properties of their host outflow (see Fig. 13). This has important implications for enhancing the applicability of the FB22 model as a sub-grid prescription across the diverse range of conditions typically encountered in cosmological simulations – one of the primary motivations for developing such models.
@@
We hope that this study will be a stepping stone in our effort to bridge the gap between idealized simulations and the complex, multiphase nature of real galactic winds. A great deal of work is still needed to generalize the behavior of cold clouds in diverse environments and constrain their collective effect on large-scale properties of their host outflows so that simulations and models can faithfully reproduce current and upcoming observations.
@@ Acknowledgements
AD acknowledges support from the Max-Planck-Gesellschaft. AD acknowledges his colleague Ayan Ghosh for helping with the cartoon illustration used in this paper. AD also expresses his gratitude to his colleague Ritali Ghosh for help with visualizations and also for several hours of useful discussions. AD also acknowledges Arif Babul, Kartick Sarkar, Evan Scannapieco Matthew Abruzzo, Volker Springel, Evan Schneider, Christoph Pfrommer, Tiago Costa, Filippo Fraternali, Greg Bryan, Rainer Weinberger, Ruediger Pakmor, Matthew Smith, Sunmyon Chon, Drummond Fielding, and Simon White for their useful comments and discussions. The authors acknowledge support from the Supercomputer Education and Research Center (SERC) at the Indian Institute of Science (IISc), Bangalore, where some of the production runs of our simulations were carried out. AD would like to express his gratitude to the super-responsive and helpful people at Paraview forum for all the technical support needed for some of the visualizations. Research of AD as a graduate student, when a significant part of this paper was written, was supported by the Prime Minister’s Research Fellowship (PMRF) from the Ministry of Education (MoE), Govt. of India. AD acknowledges the support from the Max Planck Institute for Astrophysics, Garching, for accepting him as a visiting student and providing access to their Freya & Orion supercomputing clusters when this work was initiated. Several analysis scripts in our work make use of Numpy (Harris et al. 2020), Matplotlib (Hunter 2007), and Scipy (Virtanen et al. 2020) whose communities we thank for continued development and support. MG thanks the Max Planck Society for support through the MPRG.
@@ 6 Data Availability
All the relevant data associated with this article will be shared on a reasonable request to the authors. Most of the relevant plotting and visualization codes are hosted on GitHub for public access. Follow https://github.com/dutta-alankar/CCinCC85/ for cloud-crushing in spherical CC85 wind simulations and exp_box branch in https://github.com/dutta-alankar/cloud-crushing_PLUTO/tree/exp_box where reimplementation of the expanding box simulation by Gronke & Oh 2020 has been done.
@@ Appendix A Threshold size for cloud mass growth in a uniform wind
Figure 15: The temporal evolution of cold and dense gas mass (in units of initial mass) in our ‘vanilla‘ cloud-crushing simulations (in spherical coordinates) for clouds of different sizes. The background wind is uniform and plane-parallel. In the simulations, cold and dense gas is identified as gas cells with a temperature lower than $2T_{\rm cl}=8\times 10^{4}\rm\ K$. The set of constant parameters chosen for these simulations is $(\mathcal{M},\chi)=(1.496,100)$. The parameter $t_{\rm cool,mix}/t_{\rm cc}\propto R_{\rm thres}/R_{\rm cl}$ is set to different values by changing the cloud size. Time is normalized by the cloud crushing time $t_{\rm cc}=\sqrt{\chi}R_{\rm cl}/v_{\rm wind}$. This plot shows that the transition from destruction to the growth regime takes place at $t_{\rm cool,mix}/t_{\rm cc}\lesssim 6$ giving the dimensionless parameter $\delta\approx 6$ in Eq. 1.1.
We carry out cloud-crushing simulation in a uniform plane parallel wind in spherical coordinates (‘vanilla’ cloud-crushing) for different cloud sizes to numerically evaluate the dimensionless parameter $\delta$ in Eq. 1.1. Fig. 15 shows that the transition from growing clouds to destroyed clouds (and vice versa) occurs at $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}\approx R_{\rm thres}/R_{\rm cl}\approx 6$, where the symbols have their usual meaning. We have used this numerically determined threshold to choose different cloud sizes in all the simulations presented in this paper. Such a calibration is necessary because of slight differences in the cooling function and other parameters in various existing radiative cloud crushing simulations.

### v3

@@ 4.3 Caveats and Future Directions
The Chevalier & Clegg 1985 (CC85) model of an adiabatically expanding outflow makes dramatic simplifications. Observations show that the thermodynamic state of the outflow can be significantly different from a steady CC85 wind (Xu et al. 2023). For instance, Lopez et al. 2020 observe much shallower density and temperature profiles (also in free electron density profile observed by Deanne Fisher; private communication) for the X-ray bright hot gas beyond $1\ \rm kpc$ in M82 compared to the CC85 model. Such deviations can be generic and wind profiles can potentially be non-steady in various starburst-driven outflows. Our work highlights that the expansion of the wind bears non-trivial implications for the survival/growth of cold clouds in outflows. [^9] For instance, in an isothermally expanding wind, we expect a significantly reduced cloud mass growth than in an adiabatically expanding wind. From our analytic considerations, we expect $m_{\rm cl}\propto d_{\rm cl}^{1/4}$ in an isothermal wind (not validated against any simulations carried out in this paper). This is not surprising, and depending on the background profile, clouds behave significantly differently in how they mix, shred, or grow (cf. also Tan et al. 2023 for clouds in-falling in a stratified medium).
@@
An extensive study of the behaviour of cold clouds in diverse environments of astrophysical relevance under a unified framework has yet to be explored. Note that for the majority of our runs, the cooling time of the hot medium is longer than the runtime. However, in some runs (with the largest clouds) the runtime is comparable with the wind cooling time, and we would expect thermal instabilities to occur. Given that we ignore additional heating sources (like the UV background radiation and strong stellar radiation from starburst regions irradiating the low density wind) as well as the back-reaction of the cold clouds on the wind (Schneider et al. 2018; Fielding & Bryan 2022), a self-consistent investigation of the hot gas cooling is beyond the scope of this work. Furthermore, the assumptions that go into the CC85 model, such as spherical symmetry and continuous energy and mass injection, are extremely simplistic. Supernovae can be clustered or spread out in the galactic disk, and the injection of mass and energy by them can be in the form of bursty episodes (Schneider et al. 2020; Schneider & Mao 2024; Mohapatra et al. 2025), for which the outflow does not attain a smooth CC85 profile (Sharma et al. 2014). All of these effects can add additional complications, including driving a significant amount of turbulence in the wind (Ghosh et al. 2025) and transport of metals (Sharda et al. 2024) that can potentially alter how the individual clouds or their population evolve in outflows. Therefore, different models of supernovae feedback (Warren et al. 2025) and turbulence give rise to a great deal of spatial and temporal variations in the background conditions (Mohapatra & Sharma 2019; Mohapatra et al. 2020) that different cold clouds encounter. These must be carefully taken into account before robust comparison can be made with observations and simulations (both large-scale outflow simulations of isolated galaxies; Schneider & Robertson 2017; Schneider et al. 2018, 2020; Kim & Ostriker 2018; Vijayan et al. 2018, 2020; Steinwandel et al. 2024; Wang et al. 2024; Li et al. 2025; Vijayan et al. 2025 or cosmological simulations of galaxy formation; Davé et al. 2020; Ramesh et al. 2025) of cold clouds and their host galactic winds. Although not directly related to the idealised simulations discussed here, similar caveats must be considered while testing the applicability of different sub-grid models of cold gas in cosmological simulations (Weinberger & Hernquist 2023; Smith et al. 2024a, b; Bennett et al. 2024; Butsky et al. 2024; Das et al. 2024) under various kinematic and thermodynamic conditions.
@@
In addition, the effects of thermal conduction and magnetic fields can be significant in determining the small-scale structure of cold clouds. Several recent works have demonstrated that magnetic fields suppress mixing-induced hydrodynamic instabilities, reducing the rate at which the cold mass grows in cloud-crushing simulations (McCourt et al. 2015; Kooij et al. 2021) – a picture also supported by recent idealised simulations of radiative turbulent mixing layers with magnetic fields by Das & Gronke 2024 and wind-blown bubble simulations by Lancaster et al. 2024 where suppression of hydrodynamic mixing gets manifested as a reduction in the fractal structure of turbulent boundary layers. Moreover, magnetic fields give rise to additional drag on the cloud due to draping by magnetic fields (Dursi & Pfrommer 2008) which is very sensitive to not just the magnetic field strength but more importantly to the orientation of the magnetic fields (Hidalgo-Pineda et al. 2024; Kaul et al. 2025) with respect to the direction of motion of the clouds or the direction of spatial gradients in the background medium. Strength and orientation of magnetic fields might also be responsible for the generation of cold filamentary structures observed in galactic outflows (like M82) that bear striking similarity with solar prominences. In systems similar to cloud-crushing, magnetic fields have also been found to modulate the properties of ram-pressure stripped tails in jellyfish galaxies (Müller et al. 2021). Therefore, magnetic fields and their associated effects can give rise to slow movement of long-lived clouds, leading to higher cold-mass seeding at any position in the bulk of the outflow, albeit at a lower rate. Furthermore, a distribution of multiple clouds (common in the CGM; cf. Liang & Remming 2020; Bisht et al. 2025; Hummels et al. 2024; Tan & Fielding 2024) can interact with each other (Nguyen et al. 2024) and can also change the bulk properties of the wind by inducing precipitation (owing to thermal instability; Thompson et al. 2016; Schneider et al. 2018; Dannen et al. 2020; Lochhaas et al. 2021). Such effects can give rise to multiphase outflows with significantly different kinematic and thermodynamic properties, not considered in our work. Careful consideration of such effects and their relative importance and diverse outflow environments should be considered in the future. This would enable simulations to make realistic predictions of cloud properties in different regimes and bridge the discrepancies between predictions from small-scale simulations and results of large-scale simulations of individual galaxies or cosmological simulations.
@@ 5 Summary & Conclusion
The formation and survival of cold clouds in galactic outflows are crucial for explaining the ubiquity of cold gas in the CGM. Observations suggest that cold clouds are accelerated farther from the galactic disk by powerful bursts of supernovae. The gas in an outflow is routinely found to have an expansion in the wind away from the centre of the galactic disk. In this work, we simulate the evolution of cold clouds seeded in an expanding wind, which is modelled by a steady-state Chevalier & Clegg 1985 (CC85) profile. The adiabatically expanding outflow shreds, stretches, and mixes the cold clouds seeded in the wind. In general, we find that
@@
    Compared to the indefinite growth of cold mass seen in ’vanilla’ cloud-crushing simulations (Brüggen & Scannapieco 2016; Armillotta et al. 2017; Gronke & Oh 2018; Li et al. 2020; Sparre et al. 2020; Kanjilal et al. 2021; Abruzzo et al. 2023), clouds with the same initial condition but exposed to a steady adiabatically expanding (CC85) wind, instead of a uniform wind, undergo significantly lower cold mass growth (can be lower by almost an order of magnitude; see Fig. 5).
@@
    We find that the expansion associated with the CC85 wind results in clouds embedded in this wind to spread in directions orthogonal to the wind. Once entrained, the cloud expansion is proportional to the distance travelled by the cloud $d_{\rm cl}$ (see Fig. 10), and the cold gas mass $m_{\rm cl}$ grows proportional to the square of the distance they travel, i.e., $m_{\rm cl}\propto d_{\rm cl}^{2}$ (for entrained clouds; see Fig. 11). This also explains the suppression of cold mass growth compared to the uniform wind ‘vanilla’ cloud-crushing, where $m_{\rm cl}\propto d_{\rm cl}^{3}$ – a significantly steeper cold mass growth with distance.
@@
    We compare our simulation results with the prediction of the phenomenological model of single-cloud evolution in galactic outflows proposed by Fielding & Bryan 2022 (FB22; Section 3 in their paper). Upon comparing cloud mass growth with time with our simulations using the parameters chosen for their model (Section 3.2 in their paper), we are unable to reproduce the suppression of cold mass growth in a CC85 wind compared to their ‘vanilla’ cloud-crushing counterparts (uniform wind) for all our simulation parameters. The FB22 model is very sensitive to the fractal dimension of the radiatively cooling turbulent boundary layer, which can be tweaked to match our simulation result (see Fig. 13). This highlights that the treatment of an individual cloud evolution in FB22 (Section 3.1 in their paper) is generic and can potentially be improved to match simulation results in the future. However, this is not robust across different simulation parameters and further work is needed towards developing a sound physical interpretation of any specific choice of single-cloud evolution models. FB22 also developed a general framework for the profile of a multiphase outflow (equations in Fig. 3 & Eq. 17 in FB22) with different energy and mass source terms in the wind. These source terms can come from any model of single-cloud evolution in galactic outflows. In particular, we introduce source terms to account for the cloud-wind interaction from the phenomenological cloud-crushing model that we discussed earlier. Using the general multiphase wind framework from FB22, we find that changes in the behaviour of individual clouds in the wind lead to substantial changes in the global properties of their host outflow (see Fig. 14). This has important implications for enhancing the applicability of the FB22 model as a sub-grid prescription across the diverse range of conditions typically encountered in cosmological simulations – one of the primary motivations for developing such models.
@@
We hope that this study will be a stepping stone in our effort to bridge the gap between idealised simulations and the complex, multiphase nature of real galactic winds. A great deal of work is still needed to generalise the behaviour of cold clouds in diverse environments and constrain their collective effect on large-scale properties of their host outflows so that simulations and models can faithfully reproduce current and upcoming observations.
@@ Acknowledgements
AD acknowledges support from the Max-Planck-Gesellschaft. AD acknowledges his colleague Ayan Ghosh for helping with the cartoon illustration used in this paper. AD also expresses his gratitude to his colleague Ritali Ghosh for help with visualisations and also for several hours of useful discussions. AD also acknowledges Arif Babul, Kartick Sarkar, Evan Scannapieco, Matthew Abruzzo, Volker Springel, Evan Schneider, Christoph Pfrommer, Tiago Costa, Filippo Fraternali, Greg Bryan, Rainer Weinberger, Ruediger Pakmor, Matthew Smith, Sunmyon Chon, Drummond Fielding, Simon White, and our anonymous referee for their useful comments and discussions. The authors acknowledge support from the Supercomputer Education and Research centre (SERC) at the Indian Institute of Science (IISc), Bangalore, where some of the production runs of our simulations were carried out. AD would like to express his gratitude to the super-responsive and helpful people at Paraview forum for all the technical support needed for some of the visualisations and simulation data analysis. Research of AD as a graduate student, when this work initiated, was supported by the Prime Minister’s Research Fellowship (PMRF) from the Ministry of Education (MoE), Govt. of India. AD acknowledges the support from the Max Planck Institute for Astrophysics, Garching, for accepting him as a visiting student and providing access to their Freya & Orion supercomputing clusters when this work was initiated. Several analysis scripts in our work make use of Numpy (Harris et al. 2020), Matplotlib (Hunter 2007), and Scipy (Virtanen et al. 2020) whose communities we thank for continued development and support. MG thanks the Max Planck Society for support through the Max Planck Research Group, and the European Union for support through ERC-2024-STG 101165038 (ReMMU).
@@ 6 Data Availability
All the relevant data associated with this article will be shared on a reasonable request to the authors. Most of the relevant analysis, plotting and visualisation codes are hosted on GitHub for public access. Follow https://github.com/dutta-alankar/CCinCC85/ for the cloud-crushing simulation setup in a spherical CC85 wind and also the exp_box branch in https://github.com/dutta-alankar/cloud-crushing_PLUTO/tree/exp_box, where a re-implementation of the GO20 expanding box setup has been hosted. Several videos accompanying this work are hosted in the following playlist: https://www.youtube.com/playlist?list=PL35YR13Ha_K1P721H_ub_13fmsfp6EAA2 and follow https://github.com/dutta-alankar/FB22-model-comparison.git for other relevant codes related to FB22 model comparison.
@@ Appendix A Threshold size for cloud mass growth in a uniform wind
Figure 16: The temporal evolution of cold and dense gas mass (in units of initial mass) in our ‘vanilla‘ cloud-crushing simulations (in spherical coordinates) for clouds of different sizes. The background wind is uniform and plane-parallel. In the simulations, cold and dense gas is identified as gas cells with a temperature lower than $2T_{\rm cl}=8\times 10^{4}\rm\ K$. The set of constant parameters chosen for these simulations is $(\mathcal{M},\chi)=(1.496,100)$. The parameter $t_{\rm cool,mix}/t_{\rm cc}\propto R_{\rm thres}/R_{\rm cl}$ is set to different values by changing the cloud size. Time is normalised by the cloud-crushing time $t_{\rm cc}=\sqrt{\chi}R_{\rm cl}/v_{\rm wind}$. This plot shows that the transition from destruction to the growth regime takes place at $t_{\rm cool,mix}/t_{\rm cc}\lesssim 6$ giving the dimensionless parameter $\delta\approx 6$ in Eq. 1.1.
We carry out cloud-crushing simulation in a uniform plane parallel wind in spherical coordinates (‘vanilla’ cloud-crushing) for different cloud sizes to numerically evaluate the dimensionless parameter $\delta$ in Eq. 1.1. Fig. 16 shows that the transition from growing clouds to destroyed clouds (and vice versa) occurs at $t_{\rm cool,mix}/t_{\rm cc}|_{\rm ini}\approx R_{\rm thres}/R_{\rm cl}\approx 6$, where the symbols have their usual meaning. We have used this numerically determined threshold to choose different cloud sizes in all the simulations presented in this paper. Such a calibration is necessary because of slight differences in the cooling function and other parameters in various existing radiative cloud-crushing simulations.